# Self-Generated Critiques Boost Reward Modeling for Language Models

### v1

@@ 3.2 Critique-augmented Reward Model Training
Since $p^{*}(\cdot\mid y,x)$ stands for the oracle distribution for critiques and is often not intractable, we aim to leverage the critic generation model $g_{\phi}$ to generate the approximate distribution $q_{\phi}$ as
$p\left(y^{+}\succ y^{-}\mid x\right)$ $=\sum_{z^{+},z^{-}}q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}.$ (4)
Then, by applying the Jensen’s Inequality, the training objective can be expressed as
$\mathcal{L}$ $=-\log p\left(y^{+}\succ y^{-}\mid x\right)$ (5)
$=-\log\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$
$\geq\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[-\log\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$

### v2

@@ 3.2 Critique-augmented Reward Model Training
Since $p^{*}(\cdot\mid y,x)$ stands for the oracle distribution for critiques and is often not intractable, we aim to leverage the critic generation model $g_{\phi}$ to generate the approximate distribution $q_{\phi}$ by applying the Jensen’s Inequality as
$\log p\left(y^{+}\succ y^{-}\mid x\right)$ $=\log\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$ (4)
$\geq\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[\log\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$
Then, instead of directly optimizing the negative log-likelihood, the training objective can be expressed as
$\mathcal{L}$ $=\mathbb{E}_{q_{\phi}\left(z^{+}\mid y^{+},x\right),q_{\phi}\left(z^{-}\mid y^{-},x\right)}\left[-\log\frac{p\left(y^{+}\succ y^{-},z^{+},z^{-}\mid x\right)}{q_{\phi}\left(z^{+}\mid y^{+},x\right)q_{\phi}\left(z^{-}\mid y^{-},x\right)}\right]$ (5)